# CodePlan: Unlocking Reasoning Potential in Large Language Models by Scaling Code-form Planning

### v1

@@ B.2 Efficiency Analysis
Table 10: Efficiency of CodePlan compared with vanilla training.
@@ B.3 Case Study
We present one example for each type of task in Table 1112,  1314, respectively. Note that the example for the instruction-following task has been shown in Figure 1 in the main paper. As exemplified, the code-form plan clearly outlines the step-by-step reasoning process:
@@
  • Decision-Making (Table 14): In this task set in a simulated environment, the agent must execute a sequence of actions to accomplish the given goal. While PT (Vanilla) struggles to maintain a coherent plan of action, CodePlan’s code-form plan clearly specifies the high-level steps required, such as locating and obtaining the target object, cooling it, and placing it in the designated location. This explicit planning stage allows CodePlan to systematically break down the complex task and execute the necessary actions to achieve the desired outcome.

### v2

@@ B.2 Efficiency Analysis
Table 10: Efficiency of CodePlan compared with vanilla training.
@@ B.3 Case Study
We present one example for each type of task in Table 11121314, respectively. Note that the example for the instruction-following task has been shown in Figure 1 in the main paper. As exemplified, the code-form plan clearly outlines the step-by-step reasoning process:
@@
  • Decision-Making (Table 14): In this task set in a simulated environment, the agent must execute a sequence of actions to accomplish the given goal. While the vanilla-trained LM struggles to maintain a coherent plan of action, CodePlan’s code-form plan clearly specifies the high-level steps required, such as locating and obtaining the target object, cooling it, and placing it in the designated location. This explicit planning stage allows CodePlan to systematically break down the complex task and execute the necessary actions to achieve the desired outcome.